### v4

# 4.1. Degenerate and non-degenerate components Let us use the notation in (3.3) for the Littlewood-Paley partition of unity. Typically, runs over the set of integers, but also we are going to consider the functions where runs over the set of integers and is fixed, and then In terms of the projective co-ordinates in (3.1), in view of the homogeneity of the phase function as defined in (4.2), we can write when In such a case we are going to simplify the notation by writing

4.1. Degenerate and non-degenerate components
Let us use the notation in (3.3) for the Littlewood-Paley partition of unity. Typically, runs over the set of integers, but also we are going to consider the functions where runs over the set of integers and is fixed, and then
In terms of the projective co-ordinates in (3.1), in view of the homogeneity of the phase function as defined in (4.2), we can write when In such a case we are going to simplify the notation by writing
Using the Littlewood-Paley decomposition for the Fourier integral operator we have
where the inequality with is justified because the symbol is supported on Now, we decompose with
where, as above,
and
where Here, is the “curvature” associated to the real-valued phase function Since sometimes we also write
for simplicity.
Note then that is given by
4.1.1. Symbol classes in projective co-ordinates
The aim of this subsection is to analyse the behaviour of symbols in the frequency variable with respect to the projective co-ordinates defined via
Although and are equal in value, the radial derivative keeps fixed, and the vertical derivative keeps fixed. In the next lemma we analyse the derivatives of a symbol in terms of the radial and of the angular variables. We note that and we write when and also for a row vector.
Lemma 4.1.
Let us assume that the domain of contains an open neighborhood of the region where is fixed. Assuming smooth in we have that
for some family of coefficients . Also,
on
Proof.
Let us prove (4.7). When
the chain rule gets
When observe that
Mathematical induction on
implies that can be written in terms of some coefficients as
For the proof of (4.10)
observe that assuming this identity valid for we have
With we apply (4.9) to deduce that
for some family of coefficients The proof of (4.7) is complete. For the proof of (4.8) note that for any
which verifies (4.8) when Assuming (4.8) with fixed, namely, that note that for any
from where the mathematical induction proves for all
For let be the class of symbols that consists of all functions which are Borel measurable in supported in the cone bundle
and satisfying the Hörmander type estimates
Lemma 4.2.
Let and let Let Consider the symbol
supported in the region
Then, we have the following symbol estimates
for in the support of
Proof.
Using that
satisfies the estimates in (4.12) and the identity (4.7) we have that
Observing that, on the region we have proved that
By the Leibniz rule we have that
In view of (4.8),
we can estimate
Since the sequence belongs uniformly in to we have that
Now, let us analyse the behaviour of the cutt-off at the “curvature” term Observe that for any
Then, we have the equalities
There is such that
Consequently,
Thus, the mathematical induction gives the estimate
We have proved the estimates
and
Note that when we recover the information that has order Indeed, let Since note that the order of with respect to is If we replace the analysis above with the symbol instead of then takes the place of and we get the estimate
Therefore we have proved that
The estimates for the derivatives of in and in above, the inequality in (4.20) and the Leibniz rule (4.16) conclude the proof of (4.14).
4.4. Control of the error operator
In this subsection we prove that the error operator is bounded on The idea behind the proof of this fact can be understood from the calculus of Fourier integral operators. The operator is essentially a Fourier integral operator with the same phase function that in the sense that its phase evaluated at the stationary point is see [46, Page 19, Eq. (16)]. So, intuitively, if would have had the form in (1.1), namely, one could compute an asymptotic expansion for the symbol of as follows
and satisfying the asymptotic expansion, see e.g. [27, Page 17],
and the symbols satisfying the estimates (under suitable conditions on the phase )
for any
Below we will prove that essentially and have the same principal symbol and then the order of the error operator should be, at least heuristically, the order of the ’s when that is This order is good enough to deduce the -boundedness of However, since the definition of also involves integration with respect to the structure of this operator is not as the one in (1.1), and one has to use the stationary phase method to prove that each dyadic component of the error has a kernel with a -norm satisfying uniformly in a decay proportional to Then, if proved this we are allowed to use Proposition 6.1 in [46, Page 17] to deduce the boundedness of even under the presence of the oscillating term
First, let us calculate the kernel of the error operator Note that
Consequently,
On the other hand, the kernel of is given by
In order to prove the boundedness of it suffices to prove the following estimate on each dyadic component
However, the integral inside of this -norm, can be written as follows
So, let us simplify the notation above by defining
and
Observe that in view of the Euler homogeneity identity
we can re-write the phase of the integral as follows
and then
The analysis above shows that in order to prove the boundedness of it suffices to prove the following estimate
To do so, we will estimate the integral on the set of points that contribute more information to the integrand. This set of points is determined by the region:
By following the argument in [46, Page 19] one can make the approximation
where the error term has derivatives satisfying the estimate
Denoting we have that
Here, the main integral to study is Indeed, arguing as in [46, Page 17] one can demonstrate that is a smoothing symbol. Indeed, one can estimate this integral over the region by inserting the cut-off in the integrand of Since on and in view of the identity
one has that
Repeated integration by parts in the -variable provides the factor but the differentiation of the term gives the factor which shows that this portion of the integral is for any
Note that
where
and
To evaluate we compute the integral with respect to using the distributional identity
as follows
Note that the integration with respect to is over the region and then we can estimate this integral by introducing a cutt-off
In accordance with the principle of stationary phase, we now look at where the phase is stationary in From the homogeneity of the phase function we have that
Hence
So, the only stationary points occurs when and the Hessian at this stationary point is given by
Consequently, on the support of we have the Taylor series expansion
Similarly, in view of the Taylor expansion
we can write
In view of the stationary phase principle
Consequently,
Because captures the first term of the Taylor expansion of the symbol around we have proved that is essentially a symbol of order Moreover, the factor in shows that
In consequence, the difference
has order where Now, what we want to prove is that satisfies the inequality order Since
since is a smoothing symbol the order corresponds to the order of the symbol However, in view of (4.56), one has the estimates
Since also contains the term the amplitude satisfies the estimates
The previous inequality can be written as
where
In this way, the previous estimate says that, omitting the errors is essentially a symbol of order
Now, we have prepared the way in order to use Proposition 6.1 of [46] for deducing the estimate (4.55). Indeed, our estimate in (4.55) is the analog of (19) in [46, Page 19] which is a consequence of [46, Proposition 6.1]. For completeness, we present such a proposition as follows.
Lemma 4.13.
If is a symbol satisfying estimates of the type
when then, we have the kernel estimate
So, note that for the validity of (4.55)
we have used Lemma 4.13 applied to The proof of the -boundedness of is complete.
Proof of the main Theorem 3.1.
With the notation in Subsection 4.1 the operator has been decomposed into its degenerate and non-degenerate components, and respectively. In Subsection 4.2 in Proposition 4.4 we have proved that the degenerate operator is bounded on As for the non-degenerate operator we have constructed in Subsubsection 4.3.1 a Fourier integral operator which is an average operator, of the same order that and a pseudo-differential operator of order zero. We have proved, that the factorisation approach introduced in Tao [46] also operates in our case, namely, that is bounded on (see Proposition 4.9) is of weak (1,1) type in view of the standard Calderón-Zygmund theory, and that the error operator is bounded on see Subsection 4.4. Since then is of weak (1,1) type. The proof of Theorem 3.1 is complete.